OpenStudy (anonymous):

Find the equation of the line through the point (12,15) that is parallel to 8x+3y= 8

OpenStudy (anonymous):

8x+3y=8 8x=-3y+8 x=(-3/8)y+1 Hence its slop is -3/8, and thus, so its the slope of your line, since they're parallel. So equation of line is y=(-3/8)x+c But we know also that it passes through (12,15). Thus 15=(-3/8)(12)+c 15 = -4.5+c c=15+4.5 = 19.5 = 39/2 Hence equation of line is y=(-3/8)x+39/2 8y = -3x+156 8y+3x=156

OpenStudy (anonymous):

so the answer is y=(-3/8)x+39/2)